12ft^2 to square inches

Answer by jim_thompson5910(35256)

12 inches = 1 foot


(12 inches)*(12 inches) = (12 inches)*(1 foot)


(12 inches)*(12 inches) = (1 foot)*(1 foot)


144 square inches = 1 square foot


1 square foot = 144 square inches


12*1 square foot = 12*144 square inches


12 square feet = 1728 square inches